OpenStudy (anonymous):

A line segment with endpoints U negative two comma six and V negative five comma negative three is translated according to the rule x minus two comma y plus two. Which of the following is a coordinate of the resulting line segment U prime V prime?

OpenStudy (anonymous):

Choices are: V prime negative seven comma negative five V prime negative three comma negative five U prime negative four comma eight U prime zero comma four

OpenStudy (anonymous):

@ganeshie8 @phi

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

@texaschic101 @whpalmer4

ganeshie8 (ganeshie8):

(x, y) -->(x-2, y+2)

ganeshie8 (ganeshie8):

U(-2, 6) --> U'(?, ?)

OpenStudy (anonymous):

...

ganeshie8 (ganeshie8):

just add "-2" to x coordinate, and add "6" to y coordinate

ganeshie8 (ganeshie8):

U(-2, 6) --> U'(-2-2, 6+2)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

U'(4,8)?

ganeshie8 (ganeshie8):

whats -2-2 ?

OpenStudy (anonymous):

4.

ganeshie8 (ganeshie8):

-2-2 = -4

ganeshie8 (ganeshie8):

its U'(-4, 8)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-2 x -2 = 4.

ganeshie8 (ganeshie8):

yes, thats when we multiply, but when we add -2-2 = -4

OpenStudy (anonymous):

Oh.

ganeshie8 (ganeshie8):

-2*-2 = 4 -2 - 2 = -4 -2 + 2 = 0

OpenStudy (anonymous):

But you didn't do an addition sign. :/

ganeshie8 (ganeshie8):

actually to be precise, it should be writte as below : (-2) + (- 2) = -4